# Survey in Characterization of Semantic Change

### v2

@@ 3 Corpora for Semantic Change Analysis
Corpora for semantic change could be selected based on differences in time (diachronic studies), domain (academic fields Schlechtweg et al., (2019)), culture (social media Carpuat et al., (2013); Shoemark et al., (2019), countries), language Takamura et al., (2017); Uban et al., (2021), and so on, as soon as the pair of corpora ($t_{1},t_{2}$) taken at different contexts have in common a sufficient amount of symbols for comparison.
In diachronic studies, some studies set the problem as two different data distributions with a meaningful gap between the time interval. Some methods divide the corpora into discrete bins or time windows, striking a balance between the size of the window and the quantity of data it contains. In contrast, others create no bin and proceed with a continuous representation of change where the timestamp associated with the text is directly provided to the model Hu et al., (2019); Rosenfeld and Erk, (2018); Wang et al., (2021). These corpora were designed for specific purposes. For example, the first two types of corpus rely on a discretization of word meanings in a comparative perspective, while the third one allows for an evolutionary view of the changes.
With respect to characterization, the collected papers show an increasing interest on diachronic corpora and many of them propose handcrafted rules for automatically annotating the categories of changes Tang et al., (2013). Few works rely on human annotation Schlechtweg et al., (2017). While these approaches have advantages and drawbacks, from a computational perspective, the required data to train representation models is far from being large enough. Recent state-of-the-art models for NLP require training on billions to trillions of tokens He et al., (2021), but diachronic corpora generally present tokens in a scale of millions Schlechtweg et al., (2020). Given that increasing the amount of training corpora is in most cases, infeasible, the viable approach to learning semantics in this context is with better algorithms.
Even with the progress in the fields of low-resource learning like rapid adaptation Neubig and Hu, (2018), few-shot Liu et al., (2022) or in-context learning Min et al., (2022). There is still a meaningful lack on freely available corpora for semantic change studies Kutuzov et al., (2018).
@@ 4.1.1 Dimension
To detect changes along the dimension axis ($\mathcal{D}$), Sagi et al., (2009) propose a representation method ($\mathfrak{R}$) to count together words and phonemes (e.g. ‘dog’ and ‘do’) and measure the cluster density of the cosine distance on a decomposition of the term frequency matrix. They assume that Latent Semantic Analysis is appropriate to examine vectors representing the context of each occurrence of a given term in a corpus and estimate the cohesiveness by measuring the density with which these token context vectors are related to each other in a vector space. Tang et al., (2013) propose a framework for semantic change computation by assuming that semantic change is by nature a successive and diachronic development. They further distinguish between word level (i.e. the semantic of a given word) and sense level (i.e. individual sense of a word) and implement their approach based on these two notions. One essential step in their approach aims at computing the senses of a word in a given corpus and then the diachronic distribution of these senses in the corpus. Carpuat et al., (2013) focus on neologism detection in corpora and they apply classification techniques for that. Those techniques are denoted as i) Type-level Features and encompass word frequencies change, the appearance of the word in a new context, and word’s distribution over topics change, ii) Token-level Features that refers to the same kind of features but generalize at the token level. Emms and Jayapal, (2016) propose a generative conditional model to detect if a word gained a new sense over time. This is done by modeling the word sense distribution with Gibbs sampling (a kind of a Markov chain Monte Carlo algorithm for obtaining a sequence of observations). Bochkarev et al., (2022) use a neural network to predict if a word is used as a named entity. By this method, they create a temporal view of usage of a word and they can compare the occurrences of a word to detect if it gained a new usage in the corpora.
@@ 4.1.3 Orientation
Frequency metrics to detect the co-occurrences with sets of ”good-” or ”bad-” sentiment words has been widely used to infer the amelioration and pejoration of a word over time, especially for neologism Cook and Stevenson, (2010); Jatowt and Duh, (2014). Similarly, Buechel et al., (2016) use seed words in German and English to track the emotions captured in corpora. Additionally, the authors represent it in the Valence-Arousal-Dominance framework instead of following a negative and positive connotation. A limitation of the later study is that it does not track the evolution of orientation in words.
Frequency approaches are, for a long time known to be limited for the analysis of semantic of words, given their inability to capture similarity, as stated in the work of Mikolov et al., (2013), and compositionality, as pointed out by Mitchell and Lapata, (2010). For instance, documents using different words could describe the same thing. The n-gram approach, to some extent Cheng et al., (2006), improves compositionality understanding but is hindered by the statistical significance or Zipf distribution Zipf, (1949), which is outperformed by complete sentence understanding as showed by Socher et al., 2013a .
Another limitation of frequency-based methods is its inability to represent sense and a change in word usage could not be differentiated between sense broadening or narrowing. Therefore, an extra step is required to deal with polysemy. Some authors tried normalizing frequency Hamilton et al., (2016), however, as normalization developed for one corpus is hardly applicable to another corpus Brezina, (2018), this method fell into disuse Kutuzov et al., (2018). Another approach that fell into disuse is the definition of general laws for semantic change Dubossarsky et al., (2017). The use of co-occurrence tables for a the analysis of semantic change misjudged polysemy and prototypicality. Deep study of co-occurrence tables concluded that there is no statistical correlation between frequency with polysemy or semantic change.
Finally, frequency-based methods has shown low performance to detect orientation of words Socher et al., 2013b as well as metaphor and metonymy. Since the words are deeply related to the used context, the lack of compositionality could mismatch irony or negation, and could not recognize the right ”relation” between word meanings.
@@ 4.2.1 Dimension
These approaches assume that changes in the number of topics assigned to a word can indicate changes in dimension. Lau et al., (2012) introduced the idea of novelty as the proportion of usages of a given word corresponding to senses in the focus corpus and reference corpus. The goal is to use word sense induction based on topic modeling to detect if a word is used with a new meaning. Cook et al., (2013) extended the work of Lau et al., (2012) by quantifying the use of a topic and estimating that a “cultural silent” topic that becomes frequent in the focus corpus has high probability to have neologisms. For that, they introduced the concept of relevance of a topic and compute a rank of potential neologisms. Lau et al., (2014) improved the definition of novelty by adding the frequency of the word in the corpus, then they proposed to combine novelty and relevance to rank the neologisms.
@@ 4.2.2 Relation
To the best of our knowledge, no methodology for tackling conceptual change with topic-based representation has been published, while some works use it in a synchronic detection setting Heintz et al., (2013).
@@ 4.2.3 Orientation
To the best of our knowledge, no methodology approaching orientation change with topic-based representation was published. Some works in sentiment analysis literature use topic models in a workflow in order to refine the data used by a classifier to predict the perceived sentiment. However, topic models are not directly applied for the classification task Wankhade et al., (2022).
@@ 4.3.1 Dimension
Comparing the size of a graph over time can provide rich information about the semantic change of a word. Biemann, (2006) was the first to implement this task in order to detect the sense broadening of words. They created a graph of associated nouns that occur in corpora, then applied the Chinese whispers algorithm to find clusters in the graph. By comparing these clusters over time, they were able to infer whether a word had acquired a new meaning. Mitra et al., (2015) analysed the Tweets and count the word co-occurrence based on bi-grams as a way to identify potential relationships. The graph was built based on these relationships and they run a graph-clustering algorithm and Chinese Whispers to determine the set of senses. Each cluster is considered as a different sense of a word. By comparing this sets of senses over time they can infer if a word split its senses or the senses becomes broader (joined by other words). Mathew et al., (2017) improved filtering and sense selection over previous methods introduced by Mitra et al., (2015); McCarthy et al., (2004); Lau et al., (2014), to better capture sense evolution in Google Books and newspaper datasets. Tahmasebi and Risse, (2017) consider the graph of occurrences of noun phrases with length two words as a cluster of meaning, then they merge sub-graphs based on the lin similarity. They track dimension by analysing how this sub-graphs change over time. Ehmüller et al., (2020) build a co-occurrence graph, filtering with point-wise mutual information to remove non relevant edges, and then create an ego-network to later run graph-clustering algorithms. This approach allowed to build a sense tree222a tree of word usage evolution, where new senses are new branches. Similar to previous work, the authors restricted their analysis to nouns only.
@@ 4.3.3 Orientation
The orientation of words generally is inferred from similarity or proximity with respect to seed words. Many word graphs present this construction and some works Reagan et al., (2017); Barnes et al., (2021) use these graphs to infer orientation. However, to the best of our knowledge, no work used graphs to compare the orientation of words in different contexts.
@@
Wordnet, Babelnet and Conceptnet are examples of graph structures used auxiliary to many methodologies for change detection Giulianelli et al., (2023); Tang et al., (2023). However, these structures are under-explored (generally used only as a sense storage), as they lack temporal information, like timestamps for first occurrence.
@@ 4.4 Embeddings
Word embeddings have been used for a variety of tasks, including LSC identification and characterization. This is done by comparing word vectors between two moment in time, the distance between them and with respect to the neighboring vectors can indicate if the sense of the word changed Hu et al., (2019).
@@ 4.4.3 Orientation
Methods based on word embeddings have been widely used for many sentiment analysis tasks Raffel et al., (2019) and to investigate sentiment at lexicon level Ito et al., (2020). In the context of LSC characterization, Fonteyn and Manjavacas, (2021) uses the distance between the word vectors of ‘good’ and ‘bad’ as a means to measure polarity in the expression ‘to death’.
In general, word embeddings have an unstable and many times incoherent behavior, as pointed out in Hellrich and Hahn, (2016); Gladkova and Drozd, (2016); Günther et al., (2019). The reason is that vectors representing words with the same meaning could be sensible to minor variations in the corpora or in the architecture utilized for language modeling, thus different vectors can represent the same meaning in different corpora Wendlandt et al., (2018); Leszczynski et al., (2020); Mickus et al., (2020). This undesirable effect is often observed for low frequency words in the corpus, which corresponds to most words in a vocabulary Zipf, (1949), and particularly for those with polysemy.
To mitigate the stability problem, where different training routing led to very different learned vectors, authors recurred to larger embedding sizes, quantization May et al., (2019), or alignment between trained embeddings Hamilton et al., (2016). However, the vector representation is already a burden, as standard approaches to semantic change identification suffer from meaning conflation deficiency Camacho-Collados and Pilehvar, (2018).
@@
Some works tried to address conflation deficiency by adding a vector for each sense Neelakantan et al., (2014); Liu et al., (2015), which generally requires additional annotation or steps to discriminate senses (i.e., given a word, which vector will be used).
An extension of these works includes contextualized words, and create embeddings based on pre-trained language models, generally extracted from ELMO Peters et al., (2018) or BERT Devlin et al., (2019). The main difference is that instead of using the pre-trained embeddings for each word, this approach uses the outputs of the transformer’s final (contextualized) layers to capture meaning. These outputs are then clustered or compared using cosine similarity to measure the sense of the word. For instance, the representation of the word ‘mouse’ in a text will be assigned to the correct semantic region, such as ‘animal’ or ‘computer device’.
Nevertheless, while they presented an improvement to sense detection, they produce words representation that are too contextual. Thus, words with the same meaning can have very different vectors if they are in a different context, and the representations are anisotropic. The distribution in the semantic space is not uniform, leading to non-interpretable distances in this space Ethayarajh, (2019). So for diachronic studies, this methodology is not reliable as it gives many false positives Kutuzov et al., (2022).
Recent studies try to improve the isotropic orientation and sense similarity (instead of contextual similarity) through contrastive learning, using entailment Gao et al., (2021); Reimers and Gurevych, (2019) or pre-training on similarity tasks Zhou et al., (2023). With a pre-training objective of reducing the distance between similar senses in the semantic space, these works improved pre-trained textual and sentence similarity models.
Another set of approaches attempts to model the polysemy of words directly. In the characterization literature, some works used density-based embeddings (word2gauss) to track this behavior. In Moss, (2020), the author selected some properties of this embedding to measure the dimension of change. A more in general review of embeddings capabilities and their limitations is presented in Apidianaki, (2022).
@@ 4.5 Summary
Table 2 summarizes the literature survey did in the previous sections, for semantic change characterization. In the rows we add the most relevant papers, while the columns are the main properties of LSC evaluated in this survey. The reference to the papers is in the first column. The second, third, and forth columns represent the three poles of change characterization, i.e. dimension ($\mathcal{D}$), relation ($\mathcal{R}$), orientation ($\mathcal{O}$). Finally, the technology adopted to compute and/or represent the sense ($\mathfrak{R}$) is in the fifth column. Notice that an “X” in a cell indicates that the paper applied the sense computation technology (listed in the fifth column) to analyze the corresponding “pole”.

### v3

@@ 3 Corpora for Semantic Change Analysis
Corpora for semantic change could be selected based on differences in time (diachronic studies), domain (academic fields (Schlechtweg et al.,, 2019)), culture (social media (Carpuat et al.,, 2013; Shoemark et al.,, 2019), countries), language (Takamura et al.,, 2017; Uban et al.,, 2021), and so on, as soon as the pair of corpora ($t_{1},t_{2}$) taken at different contexts have in common a sufficient amount of symbols for comparison.
In diachronic studies, some studies set the problem as two different data distributions with a meaningful gap between the time interval. Some methods divide the corpora into discrete bins or time windows, striking a balance between the size of the window and the quantity of data it contains. In contrast, others create no bin and proceed with a continuous representation of change where the timestamp associated with the text is directly provided to the model (Hu et al.,, 2019; Rosenfeld and Erk,, 2018; Wang et al.,, 2021). These corpora were designed for specific purposes. For example, the first two types of corpus rely on a discretization of word meanings in a comparative perspective, while the third one allows for an evolutionary view of the changes.
With respect to characterization, the collected papers show an increasing interest on diachronic corpora and many of them propose handcrafted rules for automatically annotating the categories of changes (Tang et al.,, 2013). Few works rely on human annotation (Schlechtweg et al.,, 2017). While these approaches have advantages and drawbacks, from a computational perspective, the required data to train representation models is far from being large enough. Recent state-of-the-art models for NLP require training on billions to trillions of tokens (He et al.,, 2021), but diachronic corpora generally present tokens in a scale of millions (Schlechtweg et al.,, 2020). Given that increasing the amount of training corpora is in most cases infeasible, the viable approach to learning semantics in this context is with better algorithms.
Even with the progress in the fields of low-resource learning like rapid adaptation (Neubig and Hu,, 2018), few-shot (Liu et al.,, 2022) or in-context learning (Min et al.,, 2022). There is still a meaningful lack on freely available corpora for semantic change studies (Kutuzov et al.,, 2018).
@@ 4.1.1 Dimension
To detect changes along the dimension axis ($\mathcal{D}$), Sagi et al., (2009) propose a representation method ($\mathfrak{R}$) to count together words and phonemes (e.g. ‘dog’ and ‘do’) and measure the cluster density of the cosine distance on a decomposition of the term frequency matrix. They assume that Latent Semantic Analysis is appropriate to examine vectors representing the context of each occurrence of a given term in a corpus and estimate the cohesiveness by measuring the density with which these token context vectors are related to each other in a vector space. While they explore the methodology for taxonomic changes analysis, they state that the method in fact measures polysemy of words. Tang et al., (2013) propose a framework for semantic change computation by assuming that semantic change is by nature a successive and diachronic development. They further distinguish between word level (i.e. the semantic of a given word) and sense level (i.e. individual sense of a word) and implement their approach based on these two notions. One essential step in their approach aims at computing the senses of a word in a given corpus and then the diachronic distribution of these senses in the corpus. Carpuat et al., (2013) focus on neologism detection in corpora and they apply classification techniques for that. Those techniques are denoted as i) Type-level Features and encompass word frequencies change, the appearance of the word in a new context, and word’s distribution over topics change, ii) Token-level Features that refers to the same kind of features but generalize at the token level. Emms and Jayapal, (2016) propose a generative conditional model to detect if a word gained a new sense over time. This is done by modeling the word sense distribution with Gibbs sampling (a kind of a Markov chain Monte Carlo algorithm for obtaining a sequence of observations). Bochkarev et al., (2022) use a neural network to predict if a word is used as a named entity. By this method, they create a temporal view of usage of a word and they can compare the occurrences of a word to detect if it gained a new usage in the corpora.
@@ 4.1.3 Orientation
Frequency metrics to detect the co-occurrences with sets of ”good-” or ”bad-” sentiment words has been widely used to infer the amelioration and pejoration of a word over time, especially for neologism (Cook and Stevenson,, 2010; Jatowt and Duh,, 2014). Similarly, Buechel et al., (2016) use seed words in German and English to track the emotions captured in corpora. Additionally, the authors represent it in the Valence-Arousal-Dominance framework instead of following a negative and positive connotation. A limitation of the later study is that it does not track the evolution of orientation in words.
Frequency approaches are, for a long time known to be limited for the analysis of semantic of words, given their inability to capture similarity, as stated in the work of (Mikolov et al.,, 2013), and compositionality, as pointed out by (Mitchell and Lapata,, 2010). For instance, documents using different words could describe the same thing. The n-gram approach, to some extent (Cheng et al.,, 2006), improves compositionality understanding but is hindered by the statistical significance or Zipf distribution (Zipf,, 1949), which is outperformed by complete sentence understanding as showed by Socher et al., 2013a .
Another limitation of frequency-based methods is its inability to represent sense and a change in word usage could not be differentiated between sense broadening or narrowing. Therefore, an extra step is required to deal with polysemy. Some authors tried normalizing frequency (Hamilton et al.,, 2016), however, as normalization developed for one corpus is hardly applicable to another corpus (Brezina,, 2018), this method fell into disuse (Kutuzov et al.,, 2018). Another approach that fell into disuse is the definition of general laws for semantic change (Dubossarsky et al.,, 2017). The use of co-occurrence tables for a the analysis of semantic change misjudged polysemy and prototypicality. Deep study of co-occurrence tables concluded that there is no statistical correlation between frequency with polysemy or semantic change.
Finally, frequency-based methods has shown low performance to detect orientation of words (Socher et al., 2013b, ) as well as metaphor and metonymy. Since the words are deeply related to the used context, the lack of compositionality could mismatch irony or negation, and could not recognize the right ”relation” between word meanings.
@@ 4.2.1 Dimension
These approaches assume that changes in the number of topics assigned to a word can indicate changes in dimension. Lau et al., (2012) introduced the idea of novelty as the proportion of usages of a given word corresponding to senses in the focus corpus and reference corpus. The goal is to use word sense induction based on topic modeling to detect if a word is used with a new meaning. Cook et al., (2013) extended the work of (Lau et al.,, 2012) by quantifying the use of a topic and estimating that a “cultural silent” topic that becomes frequent in the focus corpus has high probability to have neologisms. For that, they introduced the concept of relevance of a topic and compute a rank of potential neologisms. Lau et al., (2014) improved the definition of novelty by adding the frequency of the word in the corpus, then they proposed to combine novelty and relevance to rank the neologisms.
@@ 4.2.2 Relation
To the best of our knowledge, no methodology for tackling conceptual change with topic-based representation has been published, while some works use it in a synchronic detection setting (Heintz et al.,, 2013).
@@ 4.2.3 Orientation
To the best of our knowledge, no methodology approaching orientation change with topic-based representation was published. Some works in sentiment analysis literature use topic models in a workflow in order to refine the data used by a classifier to predict the perceived sentiment. However, topic models are not directly applied for the classification task (Wankhade et al.,, 2022).
@@ 4.3.1 Dimension
Comparing the size of a graph over time can provide rich information about the semantic change of a word. Biemann, (2006) was the first to implement this task in order to detect the sense broadening of words. They created a graph of associated nouns that occur in corpora, then applied the Chinese whispers algorithm to find clusters in the graph. By comparing these clusters over time, they were able to infer whether a word had acquired a new meaning. Mitra et al., (2015) analysed the Tweets and count the word co-occurrence based on bi-grams as a way to identify potential relationships. The graph was built based on these relationships and they run a graph-clustering algorithm and Chinese Whispers to determine the set of senses. Each cluster is considered as a different sense of a word. By comparing this sets of senses over time they can infer if a word split its senses or the senses becomes broader (joined by other words). Mathew et al., (2017) improved filtering and sense selection over previous methods introduced by (Mitra et al.,, 2015; McCarthy et al.,, 2004; Lau et al.,, 2014), to better capture sense evolution in Google Books and newspaper datasets. Tahmasebi and Risse, (2017) consider the graph of occurrences of noun phrases with length two words as a cluster of meaning, then they merge sub-graphs based on the lin similarity. They track dimension by analysing how this sub-graphs change over time. Ehmüller et al., (2020) build a co-occurrence graph, filtering with point-wise mutual information to remove non relevant edges, and then create an ego-network to later run graph-clustering algorithms. This approach allowed to build a sense tree333a tree of word usage evolution, where new senses are new branches. Similar to previous work, the authors restricted their analysis to nouns only.
@@ 4.3.3 Orientation
The orientation of words generally is inferred from similarity or proximity with respect to seed words. Many word graphs present this construction and some works (Reagan et al.,, 2017; Barnes et al.,, 2021) use these graphs to infer orientation. However, to the best of our knowledge, no work used graphs to compare the orientation of words in different contexts.
@@
Wordnet, Babelnet and Conceptnet are examples of graph structures used auxiliary to many methodologies for change detection (Giulianelli et al.,, 2023; Tang et al.,, 2023). However, these structures are under-explored (generally used only as a sense storage), as they lack temporal information, like timestamps for first occurrence.
@@ 4.4 Embeddings
Word embeddings have been used for a variety of tasks, including LSC identification and characterization. This is done by comparing word vectors between two moment in time, the distance between them and with respect to the neighboring vectors can indicate if the sense of the word changed (Hu et al.,, 2019).
@@ 4.4.3 Orientation
Methods based on word embeddings have been widely used for many sentiment analysis tasks (Raffel et al.,, 2019) and to investigate sentiment at lexicon level (Ito et al.,, 2020). In the context of LSC characterization, Fonteyn and Manjavacas, (2021) uses the distance between the word vectors of ‘good’ and ‘bad’ as a means to measure polarity in the expression ‘to death’.
In general, word embeddings have an unstable and many times incoherent behavior, as pointed out by (Hellrich and Hahn,, 2016; Gladkova and Drozd,, 2016; Günther et al.,, 2019). The reason is that vectors representing words with the same meaning could be sensitive to minor variations in the corpora or in the architecture utilized for language modeling, thus different vectors can represent the same meaning in different corpora (Wendlandt et al.,, 2018; Leszczynski et al.,, 2020; Mickus et al.,, 2020). This undesirable effect is often observed for low frequency words in the corpus, which corresponds to most words in a vocabulary (Zipf,, 1949), and particularly for those with polysemy.
To mitigate the stability problem, where different training routing led to very different learned vectors, authors recurred to larger embedding sizes, quantization (May et al.,, 2019), or alignment between trained embeddings (Hamilton et al.,, 2016). However, the vector representation is already a burden, as standard approaches to semantic change identification suffer from meaning conflation deficiency (Camacho-Collados and Pilehvar,, 2018).
@@
Some works tried to address conflation deficiency by adding a vector for each sense (Neelakantan et al.,, 2014; Liu et al.,, 2015), which generally requires additional annotation or steps to discriminate senses (i.e., given a word, which vector will be used).
An extension of these works includes contextualized words, and create embeddings based on pre-trained language models, generally extracted from ELMO (Peters et al.,, 2018) or BERT (Devlin et al.,, 2019). The main difference is that instead of using the pre-trained embeddings for each word, this approach uses the outputs of the transformer’s final (contextualized) layers to capture meaning. These outputs are then clustered or compared using cosine similarity to measure the sense of the word. For instance, the representation of the word ‘mouse’ in a text will be assigned to the correct semantic region, such as ‘animal’ or ‘computer device’.
Nevertheless, while they presented an improvement to sense detection, they produce words representation that are too contextual. Thus, words with the same meaning can have very different vectors if they are in a different context, and the representations are anisotropic. The distribution in the semantic space is not uniform, leading to non-interpretable distances in this space (Ethayarajh,, 2019). So for diachronic studies, this methodology is not reliable as it gives many false positives (Kutuzov et al.,, 2022).
Recent studies try to improve the isotropic orientation and sense similarity (instead of contextual similarity) through contrastive learning, using entailment (Gao et al.,, 2021; Reimers and Gurevych,, 2019) or pre-training on similarity tasks (Zhou et al.,, 2023). With a pre-training objective of reducing the distance between similar senses in the semantic space, these works improved pre-trained textual and sentence similarity models.
Another set of approaches attempts to model the polysemy of words directly. In the characterization literature, some works used density-based embeddings (word2gauss) to track this behavior. In Moss, (2020), the author selected some properties of this embedding to measure the dimension of change. A more general review of embeddings capabilities and their limitations is presented in Apidianaki, (2022).
Current state-of-the-art methods in LSC Identification rely on fine-tuning cross-language models trained on Word Sense Disambiguation (WSD) tasks for English data. These models are then used to extract word embeddings. This approach leverages the phenomenon of transferability between languages, where English data can improve the model’s performance on tasks in other languages, such as Russian (Pivovarova and Kutuzov,, 2021) and Spanish (Zamora-Reina et al.,, 2022) tasks.
@@ 4.5 Summary
Table 2 summarizes the literature survey conducted in the previous sections, for semantic change characterization. In the rows we add the most relevant papers, while the columns are the main properties of LSC evaluated in this survey. The reference to the papers is in the first column. The second, third, and forth columns represent the three poles of change characterization, i.e. dimension ($\mathcal{D}$), relation ($\mathcal{R}$), orientation ($\mathcal{O}$). Finally, the technology adopted to compute and/or represent the sense ($\mathfrak{R}$) is in the fifth column. Notice that an “X” in a cell indicates that the paper applied the sense computation technology (listed in the fifth column) to analyze the corresponding “pole”.